### v3

# Tree networks of real-world data: analysis of efficiency and spatiotemporal scales

Abstract
This work computationally investigates an algorithm that was originally proposed for creating and searching trees of patterns under the assumption that it seeks the most efficient internal representation of the external world, namely a real-world dataset of patterns. The algorithm produces a transition from star graph to regular 1d lattice, passing through irregular scale-invariant trees, under the effect of a single control parameter. The relatively novel information-theoretic tools of the von Neumann entropy and its derivative, the specific heat, together with the recent definition of thermodynamic quantities, such as work and heat, for networks are reviewed and applied to inspect the tree structures produced by the formation protocol. We observe some consistency between the optimal thermodynamic efficiency for networks and the optimal pattern-matching efficiency, defined in terms of the trade-off between accuracy and computational cost when the former is measured at the meso-resolution scale involved in the phase transition from star graph to scale-invariant network. We adduce the hypothesis of a richer frequency content at the meso-scale of the self-organized network to explain this observation. Moreover, empirical evidence suggests that the system benefits from the vicinity to the mentioned phase transition if it is faced with higher-complexity environments, while it does not when the environment is simpler. A phase portrait of the tree-generating algorithm is provided and multiple phase transitions are identified.
Keywords Trees  $\cdot$ Networks  $\cdot$ Thermodynamics  $\cdot$ Statistical physics  $\cdot$ Machine Learning  $\cdot$ von Neumann entropy  $\cdot$ Specific heat
1 Introduction
Why complex structures emerge in a real-world environment is a fascinating question that has not yet found a fully satisfactory answer. In a recent article [1], the emergence of complex structures is devised as the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [2]. The latter is defined in terms of microscopic quantities thanks to the fairly new developments in statistical physics of fields in network theory [1, 3, 4, 5]. In living systems, natural selection is the way the information about the environment flows into the system’s genome which will then encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system. Thus, the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6]. How these traits are optimized to maximize the fitness to the environment is an open question. The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10, 11] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines. Here we challenge this hypothesis. In our work, a system seeks the best-fitness configuration given some external constraints $r$ and a high-dimensional vector collection, the dataset $\mathcal{D}$.
We take as a case study the algorithm introduced by Portegys [12] for accelerating nearest-neighbor (1-NN) pattern recognition tasks and subsequently extended to k-NN [13]. The algorithm employs a tree-formation protocol where raw data are reorganized into hierarchical tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing network. At each step, the new pattern is linked to either the current parent pattern or to one of the k-NN children of the parent pattern. The choice is based on the distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern. The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space. Figure 1 shows the three structure prototypes produced depending on the value of the control parameter $r$. For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star graph turns out. For some intermediate value of $r$ a ragged tree structure with progressively smaller distances between the patterns is produced, while for $r\rightarrow 2$ a 1d-lattice is created. The algorithm produces exclusively trees, i.e. loop-less networks, which are necessarily sparse $|E|=N-1$ where $|E|$ and $N$ is the number of links and nodes respectively. Using the metaphor introduced by Jacob for natural selection [14], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a parameterized distance-based rule.
Figure 1: Tree structures generated by the Portegys algorithm. The algorithm creates star-like networks up to 1d lattice when $r\rightarrow 2$. For intermediate values of $r$, networks resembling an arborescent tree are produced. The process is reminiscent of the transition vapor/liquid $\rightarrow$ amorphous solid $\rightarrow$ 1d lattice. Depicted networks are generated from a subset of 50 samples randomly extracted from the MNIST dataset with $r=0.05,0.95,1.50$ respectively. Classes are indicated by the 10 different colors in the color bar.
The use of a hierarchical topology, onto which a high-dimensional dataset is mapped, offers insights into the structured disorder of the data [15] on the one hand, while on a practical perspective, it limits the number of comparisons needed in matching an unknown pattern to a prototypical example. Namely, the algorithm introduced by Portegys is equipped with a pattern-searching method described in Section 3.1. Depending on the structural traits of the networks, an optimal trade-off between fast search of the best-matching pattern and accuracy in terms of node-class labels can be defined.
The Portegys algorithm has the desirable properties of being simple, explainable, and subject to a single control parameter $r$. Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning, and network science, and we can show that information-theoretical tools such as the von Neumann entropy and its derivative reflect the functional capabilities of the network. Notably, we observe some consistency between the thermodynamic efficiency and pattern-matching efficiency that we try to explain in terms of average frequency content in the self-organized network structure.
2 Theoretical framework
2.1 Statistical physics of networks
Information spreading over networked structures is governed by the Laplacian matrix [16] and its spectrum of eigenvalues encodes many relevant topological properties of the graph [17, 18]. Let us consider an undirected and unweighted network $G(V,E)$ of order $|V|=N$ with the number of links equal to $|E|=m$. The network Laplacian is defined as $\hat{L}=\hat{D}-\hat{A}$, where $\hat{D}$ and $\hat{A}$ are the diagonal matrix of the node degrees and the adjacency matrix respectively. The Laplacian operator can be used to define a density operator [5], where the Laplacian takes the role of the Hamiltonian of statistical physics [19] as in Equation 1.
Given an initial state of the network encoding the concentration of a quantity, e.g. information, in one or more nodes, $\underline{s}(0)$, the state at time $\tau$ is given by $\underline{s}(\tau)=e^{-\tau\hat{L}}\underline{s}(0)$. Where $\hat{K}=e^{-\tau\hat{L}}$ is the network propagator and is the discrete counterpart of the path integral formulation for general diffusion processes [20, 21].
As $\hat{K}$ is a positive semidefinite Hermitian matrix admits the spectral decomposition in terms of the sum of projectors $\hat{K}=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}|$. Thus, for any state at $\tau$ it holds $|\phi(\tau)>=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}||\phi(0)>$. This relation remarks that the eigenvectors of the Laplacian are akin to the wave vectors in the Fourier space defined by the spectrum of the continuum-Laplacian, while $\tau$ can be interpreted as the "wavelength" $l\sim\tau$ of the diffusion modes, thus it defines an operational resolution scale for observations [3]. Moreover, the ensemble of accessible diffusion modes at time $\tau$ can be constructed by defining a probability measure over the accessible modes decomposed by the spectral theorem. Such canonical ensemble [5, 4, 22] is encoded into the density matrix:
$\hat{\rho}(\tau)=\frac{e^{-\tau\hat{L}}}{Z}$ (1)
The elements $\rho_{ij}$ represent the normalized amount of information transferred in a diffusion process between nodes $i$ and $j$ at time $\tau$. Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [20, 22]. The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location at time $\tau$ [4]. Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
Finally, the von Neumann spectral entropy can be introduced [5] as a multi-scale descriptor of the network structure:
$S(\hat{\rho}(\tau))=-\sum_{i=1}^{N}\mu_{i}(\tau)\log\mu_{i}(\tau)$ (2)
where $\{\mu_{i}(\tau)\}_{i=1}^{N}$ is the set of $\hat{\rho}(\tau)$ eigenvalues that are related to the set of Laplacian eigenvalues through $\mu_{i}(\tau)=e^{-\tau\lambda_{i}}/Z$. The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time $\tau$ and is bounded between $[\log C_{conn},~{}\log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes. Thus the von Neumann entropy $S$ as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [22]. Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the integrated homogeneous phase where the information has propagated all over the network.
The network entropy thus counts the (logarithmic) number of informationally connected components on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/d\log\tau$ it is possible to identify the characteristic resolution scales of the network [22, 3]. For instance, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network. The latter is the typical situation of a star graph, depicted in Figure 1, or an Erdös-Rényi network. Before the peak, the informationally connected components are the single nodes, while after the peak the informationally connected components are the full network revealing a single trivial scale. On the contrary, a constant specific heat reflects the scale-invariance of the network in a resolution region where information is lost at a constant rate. Such a situation is typical of homogeneous $d$-dimensional lattices and of heterogeneous networks whose spectral density is power-law distributed in a sufficiently extended scale-interval [3]. The latter sentence makes for a definition for the scale-invariant networks [23]. The specific heat plateau height, $c_{0}$, is also related to the spectral dimension of the network $d_{s}=2c_{0}$ [3, 23].
From these considerations, a Laplacian-Renormalization group theory has been recently developed allowing for a coarse-graining of the network structures into super-nodes unraveling the mesoscale structures [3, 22]. In this context, the peaks of the specific heat, $C$, or a marked change in its slope, identify the most convenient resolution scales where to perform the coarse-graining over the informationally connected nodes, as it allows the iteration of the procedure without an excessive network reduction, and, when the network is scale-invariant, maintaining the dilation symmetry across the renormalization group procedure [19].
To conclude, we can intuitively use the diffusion parameter $\tau$ to resolve the network topological structures at different resolution scales. Thus, the necessity of analyzing complex networked systems at all levels is solved by resorting to the conjugate Fourier space, the k-space, defined by the Laplacian spectrum and the "wavelengths" of its modes.
2.2 Thermodynamics of networks
The statistical physics theory for networks introduced above allows us to view network phenomena from a new perspective and possibly to bridge the gap to network macroscopic observables. To this regard network thermodynamic quantities such as work, $W$, and heat, $Q$, have been defined in mathematical analogy to their thermodynamic counterparts [1]. Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges results in diminishing the entropy of the network. Where lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits. The entropic cost is given by:
$Q=-\frac{\delta S}{\tau}=-\frac{\log N-S}{\tau}\leq 0$ (3)
which mathematically resembles the heat dissipation in a thermodynamic process.
On the other hand, adding edges increases the information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [1, 4]. Less obvious is how, for a fixed number of edges, the degree distribution influences the flow of information. As the Laplacian governs the continuous-time edge-centric random walk, the presence of large hubs implies an enhancement of the process with faster propagation of the diffusing quantity [16, 24]. As the free energy is inversely related to the partition function, $Z=\sum_{i=1}^{N}e^{-\tau\lambda_{i}}$, the bounds of the former are given by the bounds of the latter. From $1\leq Z\leq N$, where the lower bound is given for the connected network at large $\tau\gg 1$, while the upper bound is given by a network of isolated nodes $\forall\tau$, or for a connected one at small scales $\tau=0$. Thus, the $-\log N/\tau\leq F\leq 0$ and the gain in information flow is reflected by the variation of the free energy from the case of isolated nodes:
$W=\delta F=F-F_{iso}=F+\frac{\log N}{\tau}\geq 0$ (4)
which is tantamount to the work in the canonical ensemble. Moreover, $U=W+Q\geq 0$ which implies $W\geq|Q|$. A formation process can be depicted as the trade-off between these two necessities of the networked complex system: the need to maximize the information flow, while at the same time retaining an amount of structural entropy such that the network can encode and express sufficient diversity through its traits.
Thus, to complete the analogy to thermodynamics, the trade-off between the diversity of information and information flow is captured by the efficiency [1]:
$\eta=\frac{W+Q}{W}=1-\frac{|Q|}{W}=\frac{U}{W}$ (5)
where $0\leq\eta\leq 1$ and $U\equiv-\partial_{\tau}\log Z\equiv\operatorname{Tr}[\hat{L}\hat{\rho}]$. The variational principle here exposed has been devised to macroscopically explain the sparsity and its empirical scaling law, as well as the emergence of topological features such as modularity, small-worldness, and, with partially less clear results, heterogeneity [1]. In our work, we use this variational principle to explain the structural properties of the network subclass composed of heterogeneous hierarchical trees. An additional element is that these trees are required to encode the diversity expressed in the environment they are in contact with, defined by a real-world dataset of patterns. The specific heat is used here to identify which is the resolution scale at which such a variational principle is effective for our subclass of networks during the transition to the scale-invariant regime. For scale-invariant networks, for some $\tau$-scales, $U=<\lambda>_{\tau}=c_{0}/\tau$ [3, 23], hence in this regime Equation 5 becomes $\eta=\frac{1}{\tau}\frac{c_{0}}{W}$.
3 Methods
Our results are evaluated over subsets of four different datasets: the NIST [25], the MNIST [26], the FashionMNIST [27], and the CIFAR10 [28] datasets. These datasets contain randomly collected instances of image samples out of 10 classes. It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (’sub-style patterns’) such that the sample point cloud can be ordered naturally in a network structure.
Figure 2: Dataset complexity is estimated with multiple methods for each training dataset of 2000 samples encountered in this work. In green an estimate is given computing the Lempel-Ziv complexity [29], while in orange the intrinsic dimension is estimated with the Maximum Likelihood Estimation approach [30, 31, 32]. In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each. The left axis is referred to the quantities in green and orange, while the right axis is referred to the multilayer-perceptron accuracy. We observe that the CIFAR10 dataset is the most complex one, while the NIST dataset is the least complex among the four.
The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep-learning algorithms. The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and-white images of the handwritten digits [0-9]. The FashionMNIST is a large dataset of $28\times 28$ gray-scale images of fashion products from 10 categories. The CIFAR-10 dataset contains $32\times 32$ color images belonging to 10 classes. In all our experiments a subset of 2000 samples is extracted from the train set and the test set independently. Samples from the FashionMNIST and MNIST are always binarized, while the samples from the CIFAR10 are normalized between 0 and 1 dividing each pixel value by 255, and values are rounded up to eighth decimal. In the Supplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST. The 10 classes are equally represented both in the test and in the training set. The datasets used in our experiments exhibit different levels of complexity, see Figure 2. We estimate with multiple methods the dataset complexities on the same subset of 2000 training samples used in the rest of the manuscript. The first heuristic approach is based on the test accuracy of a small multilayer perceptron (2 layers with 128 units each). In pattern recognition, the complexity of a dataset is determined by the complexity of the curvilinear hyperplanes that would be needed to optimally separate patterns from different classes. A simple but effective parameter to gauge this facet of complexity is the accuracy that a classifier (such as the MLP) can achieve in determining the true class label for a data point. Thus, a more complex dataset would result in a lower accuracy for neural networks of fixed size when compared to a less complex dataset. The second method we use is the computation of the Lempel-Ziv complexity [29] of the aggregated training set. In this case, samples need to be binarized. The third method concerns the measure of the intrinsic dimension [15, 32] of the dataset using the Maximum Likelihood Estimation approach introduced in [30] with the correction proposed in [31]. A given dataset of gray-scale images of size $pixel\times pixel$ a priori lies on a high-dimensional space of $\mathbb{R}^{pix\times pix}$, nevertheless, because it might contain a specific class of objects that account for the "world" of our problem, for instance, handwritten digits, it spans a lower dimensional manifold. Even though the definition of such a manifold is an open problem, it is possible to estimate its dimension based on the distance of the $n$-neighbouring points [15, 32]. We set the hyper-parameter $n=3$.
While in Section 4 we discuss the full range of $r$, in Section 5 we limit the discussion to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency, quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$. The Portegys algorithm we use is the extended version with $k=5$ introduced in [13] where the distance between patterns is the Euclidean distance.
Figure 3: Search method based on relative distance. In (a) the growth process is illustrated. At each step, a new pattern is linked to either the current parent pattern, $P$, or to one of its children, $C_{i}$. The parameter $r$ modulates the region of attraction of the child, indicated by the gray circle around $C_{1}$, and the one of the parent node. Points falling within the distance $r\cdot d(P,C_{i})$ are linked to the child, otherwise, an edge will be defined between the parent node $P$ and the new pattern. In (b) the search method based on relative distance is illustrated. Once a search pattern arrives at node $P$ it must choose to search either the subtree of $C_{1}$ or $C_{2}$. The choice is based on minimizing the distance to the expected patterns contained in the subtree of $C_{i}$. In the depicted case, patterns closer to the search pattern $S$ are expected to be in the subtree of $C_{2}$. Adapted from [33].
3.1 Pattern-matching efficiency
After the tree formation described in Section 1, the Portegys algorithm enters the search phase, where the best match to an unknown search pattern, $S$, is given using a first-best approach. The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list while providing a considerable speed up due to the hierarchical organization into a tree-based structure. The growth process, described in Section 1 and depicted in panel (a) of Figure 3, ensures that subtrees of a given node have on average decreasing distances for decreasing hierarchical levels. This provides the possibility of using a search technique based on relative distances. The search method is depicted in panel (b) of Figure 3 and is an iterative process that begins at the root-node pattern. Once a search pattern, $S$, arrives at a certain parent node, $P$, it must decide to continue to search for the best match into one of its children subtree, $C^{\prime}_{1}$, $C^{\prime}_{2}$ in Figure 3. The choice is taken minimizing the Portegys search-distance: $d(S,C_{i})_{search}=d(S,C_{1})-r\cdot d(P,C_{i})=d(S,C^{\prime}_{i})$, which quantifies the distance between the $S$-pattern and the potential patterns appearing in the subtree of each child $C_{i}$, collectively indicated by the gray circle $C^{\prime}_{i}$ in panel (b). Where $d(\cdot)$ indicates any metric distance equipped with semi-positivity, symmetry, and triangular inequality properties. If $d(S,C_{i})_{search}$ is less than zero then it is set to zero. In panel (b) the bold segments in red indicate the search distance to each child node and equivalently the expected distance between the unknown pattern $S$ and their expected neighborhoods of patterns $C^{\prime}_{i}$.
It is natural to define a pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
$\theta=\frac{\text{A}}{\log N_{dist}}$ (6)
the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$, where $N$ is the number of nodes. For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases, potentially decreasing the class-match accuracy. When defining the pattern-matching efficiency we are making the simplified assumption that the pattern-matching algorithm runs on a physical substrate such that visiting any node has a fixed cost in terms of energy or terms of general resources. Hence, the efficiency $\theta$ quantifies the trade-off between the cost of used resources and the gain in performance, as depicted in panel (c, f) of Figure 4 for the MNIST and CIFAR10 respectively. Note in panel (d) that the CIFAR10 dataset has a peak for accuracy which we interpret by saying that for complex datasets (see also the case of the FashionMNIST in Section S7) some network structures do possess the ability to better encode information of the "world" in their traits. We postpone the discussion of this finding to Sections 5 and 6.
Figure 4: Pattern-matching efficiency in (c, f) measured over the NIST and CIFAR10 respectively. The efficiency $\theta$ is defined as the trade-off between the classification accuracy, (a, d), and the logarithm of the number of distances computed to find the best match, (b, e). The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency $\theta$. The shaded grey area designates the standard deviation. The purple shaded area in panel (c, f) indicates a tolerance $r$-domain of optimal efficiency. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by the yellow shaded area. Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST and CIFAR10 datasets. The 10-digit classes are equally represented both in the train- and test sets.
4 Phase portrait
Figure 5: Multiple phase transitions in hierarchical tree networks. In (a) the specific heat $C$ across the range $r\in[0,2]$. At $r\sim 1.35$, indicated by the dashed white line, the hierarchical network transits from the star graph to the amorphous configuration with the emergence of scale-invariant traits evidenced by a plateau in the specific heat. The scale-invariant region shows a second phase transition from amorphous to regular 1d-lattice at $r\simeq 1.05$ indicated by the dotted white line. For $r>0.35$, the network is always scale-invariant with decreasing spectral dimension up to the transition point, above which the spectral dimension saturates to $d_{s}=1=2\cdot c_{0}$, where $c_{0}$ is the plateau height in (c). Note that at the transition point $d_{s}=4/3$. In panel (c) the entropy (blue) and the specific heat (red) for $r=\{0.25,~{}0.35,~{}0.40,~{}0.70,~{}0.90,~{}1.05,2.00\}$. Dashed blue and green lines indicate the $\tau$-scale of the first and the second peak, $\tau^{(1,2)}_{peak}$, of the specific heat respectively. In (d), (e) the modularity and the normalized diameter, $\hat{d}=d/m$, together with their respective variances (dashed lines) identify the transition points. The modularity resolves the transition involving the mesoscale structure, while the diameter resolves the transition at the larger network scale. The legend indicates the two network sizes $N=500,~{}1000$. In (f) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks quantifies the normalized (log)number of informationally connected nodes across the scale-invariant $\tau$-scales, i.e. the number of nodes included in the scale-invariant region. Note the discontinuity at the transition from star to amorphous. The dotted line shows the peak of the variance at the transition point. In (g) the entropy surface is depicted and the entropy at the first and second specific heat peaks are illustrated, see legend. In (h) the $\tau$-scales of the two specific heat peaks are illustrated with blue and green dots respectively delimiting the three dynamical phases: the segregated phase (grey), the integrated phase (light yellow) and the green region where self-similar structures are integrated at a constant-rate. In (b) exemplary hierarchical trees of $N=20$ for growing $r$ from left to right are depicted illustrating the two phase transitions at the two resolution scales. All curves in this figure are realized for networks of $N=500$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations if not specified elsewhere.
The Portegys algorithm produces tree structures as those depicted in panel (b) of Figure 5 depending on the value of the control parameter $r$. Two structural phase transitions involving two different resolution scales can be identified (see panel (a)). The transition at the largest resolution scale can be described using as an order parameter a quantity that describes the large-scale network properties: $\hat{d}=d/m$, where $d$ is the diameter of the tree and $m$ is the number of edges equal to $m=N-1$. For hierarchical trees, another choice would be the normalized number of hierarchical levels $\hat{h}=h/(N-1)$ where $h$ is the number of hierarchical levels ranging from $h\in[0,N-1]$ (see illustration in panel (b)). Both quantities describe the "volume" expansion controlled by the parameter $r$. We chose the normalized diameter $\hat{d}$ and we identify this phase transition at $r\sim 1.05$ as indicated by the peak of the variance (dotted lines) in panel (e) for the two network sizes $N=500,~{}1000$ as indicated in the legend. The network structure transits from the amorphous regime to the regular 1d-lattice. Just below the transition point, a spectral dimension of $d_{s}=4/3=2c_{0}$ characteristic of random trees [34] is measured as indicated in (a) and shown in panels (c). Above the transition $d_{s}=1$, as it is expected for a 1d-lattice. Where $c_{0}$ is the specific heat plateau height.
The second change of phase is the transition from the star graph to the amorphous configuration at $r\sim 0.35$. Here, the emergence of scale-invariant structural properties is identified by the appearance of a second peak in the specific heat and a flat plateau in between the two $C$-peaks, as shown in panels (a) and (c) of Figure 5. Before the transition, a single peak in the specific heat identifies a single trivial $\tau$-resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. The emergence of a plateau $C=c_{0}>0$ for $r>0.35$ reflects the emergence of progressively larger scale-invariant structures that for $r\rightarrow 2$ would include the full network in the thermodynamic limit $N\rightarrow\infty$. This second change of phase is the one mainly investigated in this work as the peaks in the thermodynamic and pattern-matching efficiencies occur either in the vicinity of this transition border or are included in the scale-invariant regime. This transition, conversely to what we did with the amorphous to regular-lattice transition discussed above, can’t be identified by a quantity resolving the large-scale property of the network. In fact, we observe in panel (e) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition. Nevertheless, a phase transition indeed exists and an appropriate order parameter needs to be employed to identify it.
This phase change occurs at smaller resolution scales, specifically, the mesoscales involved in the structural transition from the star graph to the scale-invariant regime (see the bifurcation in the specific heat and the related quantities in panels (a,c,g,h)). Thus, we can identify this transition using as an order parameter the modularity [35] of the tree which we show in panel (e). We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in [36]. We stress that modularity is a mesoscale descriptor while the diameter of the network is a large-scale descriptor. This motivates the choice of a mesoscale $\tau\sim 10$ employed in Section 5 that can resolve structural changes from the star graph phase to the amorphous scale-invariant phase.
In panel (f), the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks counts the normalized (log)number of informationally connected nodes across the $\tau$-scales where the entropy is consumed at a constant rate, i.e. across the specific heat plateau. In simple terms, we count the fraction of nodes included in the scale-invariant region. We observe a discontinuity at the border of the scale-invariant transition that is not observed at the border of the amorphous to regular lattice transition.
As already discussed in Section 2.1, also along the $\tau$-direction a phase transition exists [22, 3] delimited by the two peaks of $C$ in $\tau^{(1,2)}_{peak}$ and, in between the two, we can distinguish three dynamical phases (see panel (h)). For $\tau<\tau^{(1)}_{peak}$, the network is in the informationally segregated phase; while for $\tau>\tau^{(2)}_{peak}$ is in the informationally integrated phase. Only for $r>0.35$ and $\tau^{(1)}_{peak}<\tau<\tau^{(2)}_{peak}$, we observe the dynamical regime of constant-rate entropy loss characteristic of scale-invariant structures with dilation symmetry [3, 23].
5 Results
Figure 6 shows the pattern-matching efficiency defined in Equation 6 and the thermodynamic efficiency of Equation 5 for the four investigated datasets.
In panels (a)-(d) the pattern-matching efficiency, $\theta$, is depicted. The vertical red-dashed line indicates the $r$ value for the optimal trade-off, while the shaded purple region indicates a $r$-domain tolerance. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
In panels (e)-(h) the thermodynamic efficiency, $\eta$, is depicted and, for the case of the NIST and the FashionMNIST a notable consistency in the global maxima is observed with the pattern-matching efficiency $\theta$ in terms of optimal $r$. See the $r$-range spanned by the purple-shaded areas in panels (a)-(d) and the vertical dashed lines in panels (e)-(h). This consistency in the maxima of the two efficiencies suggests that thermodynamic or pattern-matching might be related to the context of this algorithm. In the Section 6 we discuss this matter.
Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star graph. Thus, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=c_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (i)-(l) for the investigated dataset. Such transition is also depicted in Figure 7 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (i)-(l) of Figure 6. See also Section S4 illustrating the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the mesoscale structural transition from star graph to amorphous scale-invariant discussed in Section 4. Such critical resolution scale is identified by the bifurcation of the specific heat peaks depicted in panels (i)-(l) of Figure 6 and in panels (a)-(c) of Figure 7. In the seminal work [1], it is indicated to compute the thermodynamic efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e. the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this provides the time-scale where the diffusion process has reached equilibrium, and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
Nevertheless, on the one hand, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$. As shown in panel (h) of Figure 5 and in panels (a)-(d) of Figure 7 where the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves are illustrated. The key difference is in the use of the thermal ensemble average which employs all the eigenvalues to compute the specific heat. Thus, it is crucial to identify which is the $\tau$-scale that resolves the phase transition from the single trivial scale of the star graph to the more complex topological traits of the amorphous network depicted in panel (b) of Figure 5. Such resolution scale is the scale we consider of significant importance for the computation of the efficiency $\eta$ across the transition and defines an "ultra-violet" cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3]. In Section S3 the efficiency, $\eta$, is computed in the range interval $\tau<10$ as is indicated by the diffusion time defined in [1] $\tau_{diff}=1/\lambda_{2}$, i.e. the red curves depicted in panels (a). Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching accuracy.
Moreover note in Figure 7 the slow down of diffusion due to the increasing "volume" of the network discussed in Section 4 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$, we remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only which shapes the degree distribution as the number of nodes and edges is fixed across the transition. We note that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
Figure 6: Thermodynamic & pattern-matching efficiency. In (a)-(d) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure). The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line. In (e)-(h) the optimal thermodynamic efficiency, $\eta$, as a function of $r$. In panels (f) and (h) a noticeable cohesion of the global maxima of the two efficiencies is observed. Vertical dashed lines indicate the global maxima of the efficiencies. Remarkably, the thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i)-(l). The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant length scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=c_{0}>0$, reflecting a roughly scale-invariant structure. Note that for datasets of higher complexity the highest efficiency, both thermodynamic and pattern-matching, tends to be posited closer to the structural phase transition to scale-invariant illustrated by the specific heat in panels (i)-(l). This is in line with the argument that states that a fairly simple dataset does not benefit from the vicinity of the phase transition. Quantities are averaged over 50 independent simulations.
Figure 7: Role of the ensemble average on the estimation of the $\tau$-scale. Panels (a)-(d) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$. $\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Figure 6. $\tau_{diff}$ is the inverse of the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$. Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (i)-(l) of Figure 6. At smaller $r$ values the equilibrium time is better reflected by $\tau^{(2)}_{peak}$. The equilibrium of the diffusion process is an increasing function of $r$. $\tau_{diff}$ curves are averaged over 50 independent simulations. The $\tau^{(1,2)}_{peak}$ curves are computed from the specific heat curves which are averaged over 50 independent simulations.
5.1 Performance at the emergence of scale-invariance
As previously discussed in Section 4, in panels (i)-(l) of Figure 6 and in panels (a)-(d) of Figure 7, the $\tau$-scale where the bifurcation shows up indicates the scale involved in the structural transition, and it is the resolution scale where the thermodynamic efficiency for some of the datasets displays consistency with the pattern-matching efficiency as a function of $r$. Nevertheless, the critical value of $r$ where such structural transition appears is not necessarily the value where the efficiencies have the global maximum.
To discuss this empirical result we turn our attention to the panels (a)-(d) and (i)-(l) in Figure 6 and panels (a)-(d) in Figure 7. For the CIFAR10 and FashionMNIST datasets there is a rough correspondence between the $r$-value where the structural transition emerges, $r\sim 0.5$ for CIFAR10 and $r\sim 0.35$ for FashionMNIST, and the $r$-range where the global maximum of the efficiencies is observed, $r\in(0.30,0.50)$ and $r\in(0.35,0.55)$ respectively for the two datasets. On the other hand, for the MNIST and NIST datasets the thermodynamic and pattern-matching efficiencies show a maximum in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.4$ and $r\sim 0.5$. The term "far away" is quantified by the purple-shaded areas in panels (a)-(d) of Figure 6. The argument also holds for the MNIST if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area. For the MNIST case, we observe that a third large peak emerges in between $\tau^{(1)}$ and $\tau^{(2)}$ (see panel (k) in Figure 6 and the entropy curves depicted in Figure S6). This effect is only mitigated for larger datasets, i.e. larger networks, as we show in Section S2 of the Supplementary. This intermediate peak is sufficiently strong to be reflected on a local maximum in the efficiency, $\eta$. Even though this local maximum in $\eta$ is not registered by the averaged pattern-matching efficiency $\theta$ in Figure 6(c), it is reflected in its standard deviation which allows us to include the $r$-range indicated by the shaded purple area, which corresponds to the $\eta$-peak-to-peak $r$-range in Figure 6(g). We stress that also in this case, we measured $\eta$ at the $\tau$ scale indicated by the bifurcation $\tau\sim 10$.
To explain the diversity in the positioning of the maximum of the efficiencies, either in the vicinity of the transition point or farther away, we turn to the argument, already present in the literature [37, 10, 38], that indicates the necessity of tuning a system, or of a system to self-tune, at the border of a phase transition only if a high level of complexity is required by the task, see Figure 2. While for the CIFAR10 and FashionMNIST, more complex datasets when compared to both the MNIST and NIST, their respective optimal value for the efficiency is posited in the proximity of the phase transition, the optimal values of $r$ for the NIST and MNIST datasets are far away from their respective transition values. The MNIST, that is the intermediate dataset in terms of complexity, is an interesting case as it shows two maxima for the thermodynamic efficiency $\eta$. While the global maximum in Figure 6(g) is far away from the critical value of $r$ illustrated by the bifurcation in panel (b) of Figure 7, the local maximum is at $r\sim 0.35-0.40$ which corresponds to the critical value of $r$ indicated by the bifurcation. This observation is in line with the hypothesis that the system takes advantage of the critical behavior of a certain resolution region only when a fairly complex task is presented to it and in line with other observations discussed in the literature [37, 10]. Thus, it stresses the importance of the dataset under analysis and its disordered distribution of points in a high-dimensional feature space.
[…]
6 Discussion
The ability of a system to survive is the capability of concentrating a flow of order from the environment - “drinking order from the environment” [39]. In other words, systems, whether living systems or machine(-learning) systems, need to learn an efficient internal representation of the complex world - or the dataset- they are posed in contact with. Such internal representation can be encoded in terms of abstract probability distributions [9, 10] or in the very network-features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [40], or in a mixed-way as in artificial neural networks where the engineer designs a complicated distribution through the definition of an architecture and the training phase fills in the parameters in order to obtain the desired function-approximation machine to achieve statistical generalization [41].
When confronted to real world environment, in order to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [42]. Moreover the same system needs to cope with cost of sending information across its structure which can cause an enhanced exploitation of resources. In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g. class-matching.
The recently developed theoretical tools of statistical physics of fields for networks [5, 4, 43, 22, 3, 1] permit to quantify the diversity in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy. As fastness of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex networked structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [1].
In our work, this variational principle is computationally employed to explain the structural properties of the network subclass composed of sparse heterogeneous hierarchical trees produced by the Portegys protocol that, similar to living systems, need to learn the diversity of the environment they are in contact with and consequently self-organize their structure. The tree-formation undergoes two phase transitions at two different scales: a meso-scale with the emergence of a approximately scale-invariant structural behaviour that progressively involves a larger scales, and a large-scale transition from the irregular scale-invariant behaviour to the regular-scale invariant 1d lattice. Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets with different level of complexity. When the dataset was more difficult, an optimal efficiency to the pattern-searching task is found to be close to the transition to scale-invariant in the parameter space. Moreover, empirical evidence suggests that the thermodynamic efficiency and the pattern-matching efficiency show some consistency, namely when we look at the global maxima of the two in the $r$-range. Even though the two processes involved possess significant differences, namely the Portegys search phase is a directed process starting from the root node, while the diffusion propagator describes an undirected process propagating from any possible node, we think that is reasonable to expect some similar behavior of the two efficiencies, as both quantities, in the end, estimate a trade-off between the flow of a process and the encoding capabilities of the system in its network structural traits when confronted to the disorder embedded in a dataset of real-world patterns.
In both cases, the "flow" is quantified by a monotonic decreasing quantity as a function of $r$. For the Portegys we have the logarithmic number of computed distances, while for the Laplacian process, we have $\delta F=F+\frac{\log N}{\tau}$, where $F=-\frac{\log Z}{\tau}$. It is easy to convince ourselves that these two quantities describe the flow in the network as the former counts the number of visited nodes across the traversing of the network by the Portegys search process, while the latter is minus the logarithm of $Z$, with $Z$ proportional to the average return probability after time $\tau$ of a continuous time edge-centric random walker [4]. They are not exactly alike as one indeed involves a directed process while the former involves an undirected process; still, they can be used as estimates of the "flow".
We have discussed the denominator of the two efficiencies: $\eta=U/W$ and $\theta=\text{Accuracy}/\log N_{dist}$. We now discuss the numerator. For what concerns the Portegys algorithm, we have observed that there are more descriptive network structures per se as evidenced by the fact that the accuracy is not necessarily a monotonic decreasing function of the control parameter $r$, but it can exhibit higher values at specific $r$-values as illustrated by the cases of the CIFAR10 and the FashionMNIST.
The matter is what makes a certain structure better suited in terms of accuracy to a given task, or equivalently in terms of a complex input-output mapping. Analogously to artificial neural networks, enlarging the number of parameters eventually results in achieving more complicated functions, even though this doesn’t imply a one to one relation between the two. In the case of a network structure the parameters are the network traits themselves, which in the Fourier space become the “network frequencies” $\lambda_{i}$, i.e. the Laplacian eigenvalues. We can measure an effective number of parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5. The internal energy $U$ is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale $\tau$, tantamount to the inverse temperature $\beta$ in statistical physics. Thus, the Gibbs average frequency content is quantified by $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$ the internal energy, where for scale-invariant networks $<\lambda>_{\tau}=c_{0}/\tau$ [3]. The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, but we can let the structure self-organize given an external control parameter $r$.
Notably, we have computed the thermodynamic efficiency at a specific resolution scale $\tau$. We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial structure, that is a single specific heat peak, to the emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an "ultra-violet" cut-off delimiting the relevant frequencies of the network structure across the transition to scale-invariance, i.e. from star graph to amorphous. The operational $\tau$-scale employed roughly identifies modules in the network as at the $r$-bifurcation point we have observed a phase transition when employing the modularity as an order parameter. While the transition point to scale-invariant permits the identification of the network resolution scale where the thermodynamic efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency. We explain this with the statement that a system benefits from the vicinity to a phase transition only when it confronts environments of higher complexity, while it does not benefit from it when the environmental complexity is lower. This is elucidated by the use of four different datasets of increasing complexity and supported by other works in the literature [37, 10, 38].
In terms of practical utility, the methods presented could in principle provide diagnostics to tree-based approaches to classification such as hierarchical clustering, even though a limitation of this study is the investigation of solely the Portegys algorithm. Given a pre-trained tree network structure on a dataset that is unknown, it is possible to qualitatively identify the complexity of the dataset and the ability of the tree structure to cope with that complexity based on the position of the thermodynamic efficiency on the phase portrait without performing any test phase. As discussed in this manuscript, if the trained network is “struggling” in the representation of the dataset, it posits itself close to the phase transition boundary at the cost of a slower pattern search. Interestingly, our analysis on hierarchical tree networks could be also applied, if further developed, to other topics in artificial intelligence. We find it particularly suited to the context of dendritic learning [44], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for biologically inspired computing. In the context of pure neuroscience, it would also be interesting to design experiments where the dendritic structure is related to the capability of the neuron to deal with complicated input-output mapping of signals. Moreover, potential applications include the topic of neuromorphic computing [45, 46, 47] currently looking for a possible solution to the ever-increasing computational power demand of recent years. In the neuromorphic paradigm computing devices are designed such that computation and memory are colocalized on the physical substrate in the attempt to emulate the brain learning capabilities in materio. The full potential of biological neural systems is likely achieved because of the interplay between their complex topological structures and the functional dynamics of a very large number of elements evolving on it [48]. In materio tree-like structures are currently physically implemented. We mention tree structures produced at the critical point of a percolation transition [49, 50, 51] or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [49, 52], thus, their capabilities could be investigated and enhanced based on this work and future developments.
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
DC: Theoretical framework, Conceptualization, Methodology, Simulations, Data curation, Software, Writing – original draft, Writing – review & editing. LS: Writing – review & editing, Conceptualization, Supervision, Funding acquisition.
Funding
This work was funded by EUs Horizon 2020, from the MSCA-ITN-2019 Innovative Training Networks program ”Materials for Neuromorphic Circuits” (MANIC) under the grant agreement No. 861153, as well as by the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen).
Acknowledgments
We are thankful to Jos van Goor for providing some of the scripts that enabled efficient grid-search across parameters and to Federico Balducci for the useful discussions.
Supplemental Data
Supplementary Material include figures that are mentioned in the main text but are omitted from it to the purpose of clarity.
Data Availability Statement
The original contributions presented in the study are included in the article/supplementary material, further inquiries can be directed to the corresponding author/s. Datasets and scripts can be found at https://github.com/CipolliniDavide/Laplacian_Trees.